OpenStudy (anonymous):

hey i'm getting ready to do a timed assignment for pre calc. could someone be prepared to help me it will only be two to three questions ?

OpenStudy (demandandsupply):

I'll have a go

OpenStudy (anonymous):

okay thanks so much:) i have a little understanding but its hard timed. it will mainly be about, arithmetic and geometric sequences,mathematical induction,parametric equations, and polar coordinates.

OpenStudy (demandandsupply):

Can try my best :)

OpenStudy (anonymous):

n=1+2(n-1)^2 solve for the 6th term

OpenStudy (demandandsupply):

So would you just substitute in 6 into the right hand side?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

geometric sequence 10 and common ratio =2

OpenStudy (demandandsupply):

Sorry, I don't know. @jdoe0001 Send help?!

OpenStudy (jdoe0001):

\(\bf \textit{looks like }a_{\color{red}{ n}}=a_1\cdot r^{{\color{red}{ r}}-1}\implies a_{\color{red}{ n}}=10\cdot 2^{{\color{red}{ n}}-1}\)

OpenStudy (jdoe0001):

though is easier to post anew, more exposure, so if we dunno someone else may, and we can revise each other